1. In some water softening operations, Mg2+ must be removed from solution along with Ca2+. What pH is required to lower TOT Mg to 10 mg/L as CaC03 by precipitation of Mg(OH)2(s)?
1
Expert's answer
2021-11-03T08:06:01-0400

10-8×10

= 10-7

Log 10-7=-7

14= pKa+pH

pH=14-7= 7